# Paying to Do Better: Games with Payments between Learning Agents

### v1

@@ 4 Single-Player Manipulations
We now turn our attention to general games with $n$ players and focus on analyzing the effects of the simplest type of manipulations: payment policies where only a single agent pays. By making sufficiently high payments, clearly, an agent could induce any pure outcome in the game. The main questions are thus not about the power of payment policies, but about when a player has the incentive to use such policies, and what are the implications of such profitable payment policies for the welfare of other players. Proofs for this section are in Appendix E.
@@
Notation: Denote the welfare of an outcome $s$ by $w(s)$ and the best-response utility of player $i$ to action profile $s$ by $u_{i}^{BR}(s)=\max_{s_{i}^{\prime}\in S_{i}}u_{i}(s_{i}^{\prime},s_{-i})$. For the following analysis, it is useful to define a notion of regret that measures the opportunity cost per step for player $i$ of playing the game where the strategy profile $x$ is played compared to the benchmark of his best action in hindsight in the (different) game where $y$ (rather than $x$) is played by the other players. To avoid confusion with standard regret (which considers the benchmark of the best response in hindsight to $x$), we call this “alternative regret.” Intuitively, it measures the extent to which a player regrets not incentivizing others to play $y$ rather than $x$.
@@
Let $x,y$ be two joint distributions on the players’ joint action space $S$. The alternative regret of player $i$ under the distribution of play $x$ compared to best responding to the alternative distribution of play $y$ is $R_{i}(x,y)=u_{i}^{BR}(y)-u_{i}(x)$.
@@
Fix a finite game $\Gamma$ and a set $A$ of regret-minimizing agents. Let the optimal welfare outcome [^9] To simplify the analysis, we consider the generic case of a unique optimal outcome. In games with equalities in utility, this can be obtained by considering infinitesimal perturbations of the utilities. of $\Gamma$ be $y^{*}$. If $y^{*}$ is a Nash equilibrium, then there is a player $i$ such that $i$ increases his own payoff and the payoffs of all the other players by using a payment policy $p_{i}$ that induces $y^{*}$ as the unique long-term outcome of the agents’ dynamics. If $y^{*}$ is not a Nash equilibrium, the same holds if there exists a player $i$ for which the welfare gap $OPT-w(x)$ is greater than $\sum_{j\neq i}R_{j}(x,y^{*})$.
@@
To assure the convergence of arbitrary no-regret agents to $y^{*}$, the payments must induce $y^{*}$ is the unique CCE, [^10] With multiple CCEs, no-regret dynamics may fail to converge even in the time average [54]. which implies that it is also the unique Nash equilibrium. One way to do so is by making $y$ an equilibrium in dominant strategies. Although weaker requirements would be sufficient for our purpose (such as inducing a dominance-solvable game), using dominant strategies simplifies the analysis. Moreover, it turns out that it leads to the same payments and the same utilities for the players in the long term as inducing a unique CCE: [62] analyzed a mechanism-design scenario where a designer adds exogenous payments to the game, and they observed that the actual cost of inducing an outcome $s$ as an equilibrium in dominant strategies — which they called the optimal $k$-implementation of $s$, $k(s)$ — is equal to the cost of inducing that outcome as a Nash equilibrium. Combining this with the fact that an equilibrium in dominant strategies is also the unique CCE, we have that, indeed, the payments needed for inducing a unique CCE or an equilibrium in dominant strategies are the same. The intuition for this result is that further payments that are intended to assure dominant strategies are in fact made only off the equilibrium path, and are thus not actually made when the game is played. A similar effect was also used in [7] in the context of providing collateral contracts to mitigate strategic risk and incentivize investments.
This intuition also carries over, to some extent, to our setting, where one of the players uses a payment policy to manipulate the agents’ dynamics. But our case is different in several ways. First, importantly, in our case there are no external funds injected into the game. Instead, the payments that a player allows his agent to make to other agents are taken from his own payoffs. Second, as we show below, the cost of inducing an outcome $s$ as an equilibrium is lower than $k(s)$, since the player who wishes to induce $s$ as the outcome of the agents’ dynamics does not make payments to his own regret-minimizing agent. Third, since in our case the learning agents do not reason about payment policies but rather learn to respond over time, some agents will play dominated actions for some time, which will lead to additional short-term costs. This is formally expressed in the following lemma.
@@
The next lemma specifies the condition under which, given a distribution $x$ to which the no-regret agents converge in expectation in $T$ rounds, there is a player who can use a payment policy to increase his payoff by pushing the dynamic to a different outcome.
@@
Fix a finite game $\Gamma$ and a set of regret-minimizing agents for the players. If there exists a player $i$ and outcome $y$ such that $w(y)-w(x)>\sum_{j\neq i}R_{j}(x,y)$, then $i$ can increase his payoff by making payments that induce $y$ as the unique outcome of the agents’ dynamics.
Intuitively, the increase in welfare must be large enough to compensate the other players for their regret for playing $x$ relative to their best responses to the alternative outcome $y$.
@@
If the dynamics without payments approach a welfare $w(x)$ lower than the welfare of some Nash equilibrium $y$, then, since in equilibrium $R_{i}(y,y)=0$, there is a player who can increase his payoff by using a payment policy that induces $y$ as the long-term outcome of the dynamics.
The lemmas above provide the basis for proving Theorem 3. See Appendix E for the proof.
From the analysis above, we observe that in many cases players can gain from using payment policies to divert the agents’ dynamics to more favorable outcomes for them. The following theorem shows that this is true for a broad class of games. For games with $PoS=1$, this is already implied by Theorem 3, but the following result is more general.
@@
The converse is not true in general: there are games with $PoA=PoS$ that are not stable. One such example is the prisoner’s dilemma game in Figure 1.
@@ 5 Two-Player Games
In our prisoner’s dilemma example in Section 1, we saw that the players had incentives to use payments between their agents, and in equilibria of the payment game, both players had higher gains than in the game without payments. We now show that these results hold quite generally for two-player interactions. The proofs for this section are in Appendix F.
@@
Every equilibrium of a two-player payment-policy game is a (possibly weak) Pareto improvement over the no-payments outcome of this game.
The intuition of the proof is that in order to have a positive payment in equilibrium, some player must benefit from making the payment while the other benefits from receiving it. If some player is worse off with the payments, then they always have a deviation that effectively cancels the payments. With more than two players, there may be players who cannot profitably counteract such payments made among others.
The next theorem addresses the question of incentives to use payments. The theorem considers games in which no-regret dynamics converge to a single CCE, and shows that despite the simplicity of learning in these games, in many cases players still have an incentive to use payment policies. The intuition for this result is that in two-player interactions, players can utilize payments to assume the role of a Stackelberg leader in the game.
@@ 6 Conclusion
Autonomous learning agents are widely and increasingly used in many online economic interactions, such as auctions and other markets. Motivated by this transition to automated systems with sophisticated AI agents, we studied the scenario where automated agents also offer monetary transfers to other agents during their dynamics. Our results show that, generally, strategic users have incentives to use payments with their learning agents, and that this can have significant implications for the overall outcomes. Our findings highlight a challenge for mechanism design in the AI era, underlining the need to better understand both agent dynamics and user incentives in automated markets.
Appendix A Further Related Work
We follow a long research tradition on learning and dynamics in games, from early work of Brown, Robinson, Blackwell, and Hannan in the 1950s [13, 20, 43, 76], through seminal work in the following decades [39, 40, 44] and continuously since [1, 15, 26, 27, 37, 46, 51, 52, 53, 54, 55, 60, 83]. See [23, 45] for an overview of the foundations of the field. The notion of regret has been central throughout this work as a tool to define learning outcomes and objectives and design algorithms to achieve them.
No-regret learning has been broadly studied in auctions, with two prominent lines of work focusing on the price of anarchy in various auction formats with regret-minimizing bidders [14, 21, 28, 46, 77, 80, 84] and on the econometric problem of inferring bidder preferences [41, 67, 68, 69]. Other work has studied automated bidding algorithms for auctions with more complex features, such as budget constraints [11, 30, 37], or learning reserve prices [22, 61, 81]. In [2, 72], regret minimization is used as a prediction model for bidder behavior. Importantly, [67, 72] provide empirical evidence from large auction datasets showing that bidder behavior in practice is consistent with no-regret learning outcomes.
Learning dynamics in first- and second-price auctions and their convergence properties have been studied in a broad range of work [12, 18, 28, 29, 36, 53, 54, 71]. In [53, 54], the authors focus on analyzing the average outcomes of regret-minimization dynamics and their implications on user incentives. These papers consider a meta-game between the users of learning agents, which is induced by the long-term outcomes of the agents’ dynamics. This is close to our model of the payment game; however, payments have a very different impact on the agents’ dynamics and they lead to different analyses and results in auctions and other games.
The question of how distributed and self-interested players can reach cooperative or efficient outcomes is, of course, a very broad topic that has been studied in various (and largely separate) fields. The classic approach to this problem is from mechanism design [66, 70] and implementation theory [47, 59]. Somewhat closer to our work are [7, 62, 88], who study how an external mechanism designer can minimize the exogenous payments he needs to make to the players in order to implement particular outcomes in the game. By contrast, we study the impact of payments that automated agents could make among themselves; our analysis does not consider a mechanism-design problem [^11] Designing payment protocols to facilitate particular types of payments between the players to optimize social welfare is an interesting research avenue, but we do not pursue it here. and there are no external payments. An additional strand of work from the reinforcement learning (RL) and artificial intelligence literature studies how distributed RL algorithms can jointly reach cooperative outcomes in sequential social dilemmas [6, 33, 50, 58, 87]. A third line of work considers an approach for reaching cooperative outcomes called program equilibrium [57, 73, 74, 85], in which each player declares a program and each program can read the commitments made in the other programs and condition actions in the game on those commitments. These models, however, are very different from our model.
Finally, the idea of using payments to incentivize desirable actions is, of course, an old one, and represents one of the cornerstones of a rich literature on contract design and principal-agent problems. We refer the interested reader to [17, 31, 56] for an overview of this field. In most work in this area, there is a “principal” who provides incentives to the agents. As a result, he faces a contract-design or an optimization problem, rather than acting as a player in the game. Notable work that is closer to our scenario includes [48], who study multi-lateral payment contracts in one-shot interactions, [42, 89], who study learning in repeated contracts, and [75], who study scenarios where a player extracts fees from other players by using binding contracts. Our setting, however, differs significantly from the contracts setting, and our focus is on when players have incentives to augment learning agents with the ability to make payments, how payments among the agents affect their learning dynamics, and what the implications for the game outcomes are.
Appendix B Prisoner’s Dilemma Games
We begin by reviewing a known result (see, e.g., [49]) that is useful for our analysis.

### v2

@@ 5 Manipulations by One Player
We now turn our attention to general games and focus on analyzing the effects of the simplest type of manipulations: payment policies where only a single agent pays. The question we aim to answer here is when does a player have an incentive to use payments with a learning agents? By making sufficiently high payments to all the other agents, clearly, an agent could induce any pure outcome in the game. The main questions are thus not about the power of payment policies, but about when are such policies also profitable, and what are the implications of such profitable payment policies for one player on the welfare of other players.
@@
Notation: Denote the welfare of an outcome $s$ by $w(s)$ and the best-response utility of player $i$ to action profile $s$ by $u_{i}^{BR}(s)=\max_{s_{i}^{\prime}\in S_{i}}u_{i}(s_{i}^{\prime},s_{-i})$. For the following analysis, it is useful to define a notion of regret that measures the opportunity cost per step for player $i$ of playing the game where the strategy profile $x$ is played compared to the benchmark of her best action in hindsight in the (different) game where $y$ (rather than $x$) is played by the other players. To avoid confusion with standard regret (which considers the benchmark of the best response in hindsight to $x$), we call this “comparative regret.” Intuitively, it measures the extent to which a player regrets not incentivizing others to play $y$ instead of $x$.
@@
Let $x,y$ be two joint distributions on the players’ joint action space $S$. The comparative regret of player $i$ under the distribution of play $x$ compared to best responding to an alternative distribution of play $y$ is $R_{i}(x,y)=u_{i}^{BR}(y)-u_{i}(x)$.
@@
Fix a finite game $\Gamma$ and a set $A$ of regret-minimizing agents. Let the optimal welfare outcome [^10] To simplify the analysis, we consider the generic case of a unique optimal outcome. In games with equalities in utility, this can be obtained by considering infinitesimal perturbations of the utilities. of $\Gamma$ be $y^{*}$. If $y^{*}$ is a Nash equilibrium, then there is a player $i$ such that $i$ increases her own payoff and the payoffs of all the other players by using a payment policy $p_{i}$ that induces $y^{*}$ as the unique long-term outcome of the agents’ dynamics. If $y^{*}$ is not a Nash equilibrium, the same holds if there exists a player $i$ for which the welfare gap $OPT-w(x)$ is greater than $\sum_{j\neq i}R_{j}(x,y^{*})$.
@@
To assure the convergence of arbitrary no-regret agents to $y^{*}$, the payments must induce $y^{*}$ is the unique CCE, [^11] With multiple CCEs, no-regret dynamics may fail to converge even in the time average [64]. which implies that it is also the unique Nash equilibrium. One way to do so is by making $y$ an equilibrium in dominant strategies. Although weaker requirements would be sufficient for our purpose (such as inducing a dominance-solvable game), using dominant strategies simplifies the analysis. Moreover, it turns out that it leads to the same payments and the same utilities for the players in the long term as inducing a unique CCE: [76] analyzed a mechanism-design scenario where a designer adds exogenous payments to the game, and they observed that the actual cost of inducing an outcome $s$ as an equilibrium in dominant strategies — which they called the optimal $k$-implementation of $s$, $k(s)$ — is equal to the cost of inducing that outcome as a Nash equilibrium. Combining this with the fact that an equilibrium in dominant strategies is also the unique CCE, we have that, indeed, the exogenous payments needed for inducing a unique CCE or an equilibrium in dominant strategies are the same. The intuition for this result is that further payments that are intended to assure dominant strategies are in fact made only off the equilibrium path, and are thus not actually made when the game is played. A similar effect was also used in [10] in the context of providing collateral contracts to mitigate strategic risk and incentivize investments.
This intuition also carries over, to some extent, to our endogenous setting, where one of the players uses a payment policy to manipulate the agents’ dynamics. But our case is different in several ways. First, importantly, there are no external funds injected into the game. Instead, payments that a player allows her agent to make to other agents are taken from her own payoffs. Second, as we show below, the cost of inducing an outcome $s$ as an equilibrium is lower than $k(s)$, since the player who wishes to induce $s$ as the outcome of the agents’ dynamics does not make payments to her own regret-minimizing agent. Third, since in our case the learning agents do not reason about payment policies but rather learn to respond over time, some agents will play dominated actions for some time, leading to additional short-term costs. This is formalized in the following lemma.
@@
Proof.
(Lemma 4). We start constructing a payment policy for the agent of player $i$ by looking at the optimal $k$-implementation payments for inducing $y$ as a dominant strategy equilibrium; we will then adjust this payment profile to fit our setting in which there are no external payments and to minimize player $i$’s cost. A key point is that the payment amount $k(y)$ includes payments to agent $i$ that an external designer would have made for inducing $y_{i}$ as the dominant strategy for agent $i$. These payments are not feasible and not needed in our case. Instead, we change the payment policy of agent $i$ in the following way: for any action $s_{i}\neq y_{i}$, agent $i$ makes a payment of $\max_{s\in S}u_{i}(s)+1$, and distributes this payment equally to the other agents. These payments make all the other actions of agent $i$ strictly dominated by the action $y_{i}$. Thus, to calculate the cost for player $i$, we need to subtract from the $k$-implementation payments the payments made to player $i$, $R_{i}(y,y)$, and add a learning-phase error term $\tilde{c}$ that results from the additional payments that agent $i$ makes when playing its dominated strategies, as well as from other times during the learning dynamics when the other agents play their dominated strategies.
Claim.
The expected learning phase cost per time step $\tilde{c}$ vanishes in the long term, that is, $\tilde{c}=o(1)$.
Proof.
$\tilde{c}$ is the expected total payment per time step that player $i$ makes when agents play strictly dominated strategies. Assume for contradiction that $\lim_{T\rightarrow\infty}\tilde{c}>0$, or that the limit does not exist. Both cases imply that the frequency at which dominated strategies are played in the agents’ dynamic is not vanishing, which contradicts the regret-minimization property of the agents. ∎
The next lemma specifies the condition under which, given a distribution $x$ to which the no-regret agents converge in expectation in $T$ rounds, there is a player who can use a payment policy to increase her payoff by pushing the dynamic to a different outcome.
@@
Fix a finite game $\Gamma$ and a set of regret-minimizing agents for the players. If there exists a player $i$ and outcome $y$ such that $w(y)-w(x)>\sum_{j\neq i}R_{j}(x,y)$, then $i$ can increase her payoff by making payments that induce $y$ as the unique outcome of the agents’ dynamics.
Intuitively, the increase in welfare must be large enough to compensate the other players for their regret for playing $x$ compared to their best responses to the alternative outcome $y$. Before proving the lemma, we have the following corollary.
@@
If the dynamics without payments approach a welfare $w(x)$ lower than the welfare of some Nash equilibrium $y$, then, since in equilibrium $R_{i}(y,y)=0$, there is a player who can increase her payoff by using a payment policy that induces $y$ as the long-term outcome of the dynamics.
Proof.
(Lemma 5). We need to specify the requirement that player $i$ will increase her own utility after making the payments that induce $y$ as the unique outcome of the dynamics. That is, $cost_{i}(y)<u_{i}(y)-u_{i}(x)$. By Lemma 4 we have that for large enough $T$, player $i$ prefers making such payments if $u_{i}(y)-k(y)+R_{i}(y,y)>u_{i}(x)$. Rewriting the left terms we have
$u_{i}^{BR}(y)-k(y)>u_{i}(x)=w(x)-\sum_{j\neq i}u_{j}(x).$
[…]
Note that all the steps in the proof of Lemma 5 hold in either direction. Thus, the lemma, and consequently the result of Theorem 3 for the case where the optimal welfare outcome is not an equilibrium, also hold in the other direction. That is, if player $i$ can increase her payoff by making payments that induce $y$ as the unique outcome, then $w(y)-w(x)>\sum_{j\neq i}R_{j}(x,y).$
The lemmas above provide the basis for proving Theorem 3.
Proof.
Suppose that $y^{*}$ is a Nash equilibrium. By definition, if some profile $s$ is a Nash equilibrium, then $R_{i}(s,s)=0$ for all $i$. Thus, by Lemma 4, any player can construct a payment policy that induces $y^{*}$ as the unique outcome of the agents’ dynamic with a cost-per-time-step of $o(1)$.
Therefore, to have a player $i$ who increases her payoff by inducing $y^{*}$ as the outcome of the dynamics, we only need to have $u_{i}(y^{*})>u_{i}(x)$. Assume for the purpose of contradiction that for all $i$, $u_{i}(x)\geq u_{i}(y^{*})$. Summing over the players, we have $\sum_{i}u_{i}(x)\geq\sum_{i}u_{i}(y^{*})$; that is, $w(x)\geq w(y^{*})$, a contradiction to the optimality of $y^{*}$. Thus, we have that for $T$ large enough such that the learning-phase costs are small, there exists a player $i$ who can increase her payoff by inducing the optimal-welfare outcome as the outcome of the agents’ dynamics. If $y^{*}$ is a non-equilibrium outcome, the conditions for $i$ increasing her payoff are given by Lemma 5, setting $y=y^{*}$. Notice also that since under this payment policy, the outcome $y^{*}$ dominates $x$ for all the agents, the utilities are a strict Pareto improvement compared to the game without payments. ∎
From the analysis above, we observe that in many cases, players can gain from using payment policies to divert the agents’ dynamics to more favorable outcomes for them. The following theorem shows that this is true for a broad class of games. For games with a price of stability (PoS) of 1, this is already implied by Theorem 3, but the following result is more general.
@@
The converse is not true in general: there are games with $PoA=PoS$ that are still not stable, so players prefer to use payments. One such example is the prisoner’s dilemma game in Figure 1.
Proof.
The result follows from combining the proof of Lemma 5 with general properties of regret-minimization dynamics. Let $\Gamma$ be a game in which the Nash equilibrium with the highest welfare, denoted $\tilde{y}$, yields a welfare of $\tilde{w}$ (where the welfare in equilibrium $\tilde{w}$ is not necessarily equal to the optimal welfare $OPT$), and there exists a CCE $x$ with welfare $w(x)<\tilde{w}$.
It is known that for any CCE of a game, there are regret-minimizing agents that converge to that equilibrium [64, 77]. In particular, there are agents that converge arbitrarily close to $x$ [^13] The proofs of this convergence result use cycles of pure actions to approximate the time average of joint distributions, and the arbitrarily close approximation is due to the density of the rational numbers. We disregard this detail in our analysis. By Lemma 5 (in particular, using Corollary 1), there exists a player who prefers using a payment policy that assures convergence to the highest welfare equilibrium $\tilde{y}$ compared to not making payments and reaching outcome $x$. Thus, we have a utility-improving deviation over the no-payments action profile and hence the game is not stable. ∎
@@ 6 Two-Player Games
In our prisoner’s dilemma example presented in Section 1, we saw that the players had incentives to use payments between their agents, and in equilibria of the payment game, both players had higher gains than in the game without payments. We now show that these phenomena apply quite generally to bilateral interactions where the players use learning agents.
@@
Every equilibrium of a two-player payment-policy game is a Pareto improvement over the no-payments outcome of this game.
Proof.
The intuition for proving this result is that in order to have any positive payment in equilibrium, some player must benefit from making the payment, while the other benefits from receiving it. If some player is worse off with the payments, then, since there are no considerations about other players, this player always has a deviation that effectively cancels the payments. With more than two players, there may be players who cannot profitably counteract such payments made among others that do not benefit them. Formally, consider a game $\Gamma$, a set $A$ of regret-minimizing learning agents for the players, and a time horizon $T$. Suppose, for the sake of contradiction, that there exists an equilibrium $p^{*}$ of the payment-policy game associated with ${\Gamma,A,T}$ where player $i$ has a lower utility than her expected utility over the $T$ rounds in the game without payments. Let $h$ denote the utility for player $i$ in the game without payments, and let $l$ denote the utility for player $i$ in the equilibrium $p^{*}$. By our assumption, we have $h>l$. Consider the following deviation from the payment policy $p^{*}$: player $i$ matches her payments to zero out all payments made by the other player and, additionally, reduces all of her other payments to zero. As a result, the two agents now observe in her dynamics the utilities of the original game $\Gamma$ without payments. Therefore, player $i$ has a utility of $h$. Since player $i$ has increased her utility from $l$ to $h$, we have a contradiction to $p^{*}$ being an equilibrium. Hence, the utilities of both players can only increase relative to the game without payments. ∎
The next theorem addresses the question of incentives to use payments in a class if games in which no-regret dynamics converge to a single CCE, including games with an equilibrium in dominant strategies. The theorem shows that despite the simplicity of learning in these games and even with dominant strategies, in many cases players still have an incentive to use payment policies. One intuition for this is that in two-player interactions, players can utilize payments to effectively assume the role of a Stackelberg leader in the game.
This result is also related to a theorem in [64] that deals with games with dominant strategies in a metagame in which users misreport their utility parameters to their learning agents. In our context of payments between the agents, the argument is different and the result is more broad, applying to a larger family of games, including all dominance-solvable games, generic fully-mixed games, and socially-concave games [40] (a sub-class of concave games in which regret-minimization dynamics converge to Nash equilibria) with a unique equilibrium.
@@
Proof.
Consider as our starting point the action profile where the players use zero payments. In the game without payments, any regret-minimization dynamic will converge to the unique coarse correlated equilibrium (which is also the unique Nash equilibrium). To show that the game is not stable for any set of regret-minimizing agents for the players, we need a utility-improving deviation from the zero-payment profile for one of the two players.
We use the fact that when the opponent does not make payments, a player has the ability to alter her policy so as to commit to playing a pure action in the long term; this can be obtained by making large payments whenever the agent plays different actions, inducing a dominant. Since for at least one of the players there is a Stackelberg-equilibrium outcome that is different from the (unique) Nash-equilibrium outcome, there exists a player who obtains a higher payoff as a Stackelberg leader in the game, so this player prefers making such a commitment. Since the other agent is also minimizing regret, this agent will eventually learn to best reply to the fixed action of the first player’s agent, and the dynamics with these payments will converge to the Stackelberg outcome of the game. Thus, we have a utility-improving deviation from the zero-payment policy, so the game is not stable. Note that this is a different result from that of Theorem 4, since the Stackelberg outcome here is not a coarse correlated equilibrium of the game $\Gamma$. ∎
@@ 7 Conclusion
Autonomous learning agents are widely and increasingly used in many online economic interactions, such as auctions and other markets. The rapid advancement in large language models and their integration into a wide array of applications further lowers barriers to communication and interaction among autonomous agents. It is not difficult to envision a future digital landscape populated by a plethora of autonomous agents, endowed with access to financial assets or fiat currencies, trading among themselves to serve the interests of their users. Motivated by this transition to automated systems with sophisticated AI agents, we studied the scenario where users let their automated agents offer monetary transfers to other agents during their dynamics.
Our results show that, in a very general sense, strategic users have incentives to use payments with their learning agents, which can have very significant implications for the overall outcomes. While focus in this work is on the general properties of payment games and their analysis in the context of auctions, studying payments between automated learners may also be of interest in a wide range of other strategic interactions, including Fisher markets, asset markets, and contracting games with multiple learning agents.
More Broadly, our findings highlight a challenge for mechanism design in the AI era, underlining the need to better understand both the dynamics of learning agents and the incentives of their users in automated markets.
Acknowledgments
We thank Noam Nisan for his valuable feedback on this project. Éva Tardos was supported in part by AFOSR grant FA9550-23-1-0410, AFOSR grant FA9550-231-0068. Joe Halpern was supported in part by AFOSR grant FA23862114029, MURI grant W911NF-19-1-0217, ARO grant W911NF-22-1-0061, NSF grant FMitF-2319186, and a grant from the Cooperative AI Foundation.
Appendix A Prisoner’s Dilemma Games
We begin by stating a simple result for the class of dominance-solvable games (see, e.g., [59]) that is useful for the following analysis. A dominance-solvable game is a game in which iterated elimination of strictly dominated strategies leads to a unique outcome. In particular, this class includes games with a strictly dominant strategy equilibrium, such as the prisoner’s dilemma game.